To buy both a new car and a new house, Tina sought two loans totalling $191,610. The simple interest rate on the first loan was 6.3%, while the simple interest rate on the second loan was 8.8%. At the end of the first year, Tina paid a combined interest payment of $16,681.66. What were the amounts of the two loans?

Answer by mananth(16949)

To buy both a new car and a new house, Tina sought two loans totalling $191,610. The simple interest rate on the first loan was 6.3%, while the simple interest rate on the second loan was 8.8%. At the end of the first year, Tina paid a combined interest payment of $16,681.66. What were the amounts of the two loans?

..
$ x ---------------- 6.3%
$y------------------ 8.8%
..
Interest on both = 16,681.66
Loan = 191,610
...
x+y = 191,610.......................1
6.3x+8.8y=16681.66*100
multiply by 10
63x+88y=16,681.66*100*10
63x+88y=16,681,660...................2
multiply (1) by -63
-63x-63y=-12,071,430
add to (2)
25y=4610230
y=184,409.2 ------8.8%
x=7,200.80 -------6.3%
m.ananth@hotmail.ca